The probability that a certain type of machine will broke down in the first month of operation is 0.1. If a firm has two such machines which are installed at the same time, find the probability that, at the end of first month, just one has broken down
Answer by ikleyn(52835)

.

    P = P(A is broken, B is not) + P(B is broken, A is not) = 0.1*(1-0.1) + 0.1*(1-0.1) = 2*0.1*0.9 = 0.18.    ANSWER